# Brain-inspired sparse training enables Transformers and LLMs to perform as fully connected

### v1

@@ Abstract
This study aims to enlarge our current knowledge on application of brain-inspired network science principles for training artificial neural networks (ANNs) with sparse connectivity. Dynamic sparse training (DST) can reduce the computational demands of training and inference in ANNs, but existing methods face difficulties in maintaining peak performance at high connectivity sparsity levels. The Cannistraci-Hebb training (CHT) is a brain-inspired method for growing connectivity in DST. CHT leverages a gradient-free, topology-driven link regrowth mechanism, which has shown to achieve ultra-sparse (1% connectivity or lower) advantage across various tasks compared to fully connected networks. Yet, CHT suffers two main drawbacks: (i) its time complexity is $\mathcal{O}(N\cdot d^{3})$- N node network size, d node degree - hence it can be efficiently applied only to ultra-sparse networks. (ii) it rigidly selects top link prediction scores, which is inappropriate for the early training epochs, when the network topology presents many unreliable connections. Here, we propose a matrix multiplication GPU-friendly approximation of the CH link predictor, which reduces the computational complexity to $\mathcal{O}(N^{3})$, enabling a fast implementation of CHT in large-scale models. Moreover, we introduce the Cannistraci-Hebb training soft rule (CHTs), which adopts a flexible strategy for sampling connections in both link removal and regrowth, balancing the exploration and exploitation of network topology. To further improve performance, we integrate CHTs with a sigmoid gradual density decay strategy, referred to as CHTss. Empirical results show that, using 1% of connections, CHTs outperforms fully connected networks in MLP architectures on visual classification tasks, compressing some networks to less than 30% nodes. Using 5% of the connections, CHTss outperforms fully connected networks in two Transformer-based machine translation tasks. Finally, using 30% of the connections, CHTss achieves superior performance compared to other dynamic sparse training methods in language modeling (LLaMA-130M) across different sparsity levels, and it surpasses the fully connected counterpart in zero-shot evaluations.
@@ 1 Introduction
Artificial neural networks (ANNs) have spurred significant advancements in various fields, such as natural language processing, computer vision, and deep reinforcement learning. The most common ANNs include several fully connected (FC) layers, which represent a significant portion of parameters in recent large language models (Touvron et al., 2023a; Zhang et al., 2022). This dense connectivity poses major challenges during both the training and the deployment phases of the models. In contrast, the brain’s neural networks inherently exhibit sparse connectivity (Drachman, 2005; Walsh, 2013). This natural design in the brain leverages sparsity, suggesting a model where the number of connections does not scale quadratically with the number of neurons. This could alleviate computational constraints, enabling more scalable network architectures.
Dynamic sparse training (DST) (Mocanu et al., 2018; Jayakumar et al., 2020; Evci et al., 2020; Yuan et al., 2021; Zhang et al., 2024b) has emerged as a promising approach to reduce the computational and memory overheads of training deep neural networks while maintaining or even enhancing model performance. Unlike Pruning methods (Han et al., 2016; Frantar & Alistarh, 2023; Zhang et al., 2024a), DST starts with an already sparse initialized network and maintains this sparsity throughout the training process. Prevalent DST methods evolve the network topology during training by iteratively removing a proportion of connections and regrowing an equivalent number of links.
@@
Apart from some detailed distinctions, the primary innovation in this field centers on the development of the regrowth criterion. A notable advancement is the gradient-free regrowth method introduced by Cannistraci-Hebb training (CHT) (Zhang et al., 2024b). This method is inspired by epitopological learning—literally meaning ‘new topology’—and is rooted in brain-inspired network science theory (Cannistraci et al., 2013; Daminelli et al., 2015; Durán et al., 2017; Cannistraci, 2018; Narula, 2017). Epitopological learning explores how learning can be implemented on complex networks by changing the shape of their connectivity structure (epitopological plasticity). CHT has demonstrated remarkable advantages in training ultra-sparse ANNs with connectivity levels of 1% or lower, often outperforming fully connected networks in various tasks. However, despite these advancements, CHT encounters two major challenges: 1) During dynamic sparse training, its rigid link selection mechanism can lead to epitopological local minima, which severely impedes the exploration of optimal network topologies. 2) The time complexity of the CHT regrowth method, CH3-L3, is $\mathcal{O}(N\cdot d^{3})$, where $N$ represents the number of nodes in the network and $d$ is the average degree. As the network becomes denser, the time complexity escalates to $\mathcal{O}(N^{4})$, rendering it impractical for large-scale and higher density models.
@@
In this article, we present the Cannistraci-Hebb Training soft rule (CHTs), which introduces several key innovations: 1) To address the issue of epitopological local minima, CHTs employs a multinomial distribution to sample link scores from removal and regrowth metrics, enabling more flexible and effective exploration of network topologies. 2) CHTs incorporates the novel substitution node-based link prediction mechanisms, reducing the computational time complexity to $\mathcal{O}(N^{3})$. This improvement makes CHTs scalable to large-scale models with higher network density. 3) CHTs leverages network science principles to initialize sparse topologies for bipartite networks with small-world properties, demonstrating superior performance compared to the traditional Erdos-Renyi model. Additionally, we propose a sigmoid gradual density decay strategy, which, when integrated with CHTs, forms an enhanced framework termed CHTss. This combined approach further optimizes the training process for sparse neural networks.
@@
Figure 1: Illustration of the CHTs process. One training iteration follows the path of (a1) >->- > (b1) >->- > (c1) >->- > (c2) >->- > (d1) >->- > (e). (a1) Network initialization with each of the sandwich layers being a bipartite small-world (BSW) network. (a2) One sample BSW network with different $\beta$ values (for $\beta$ = 0, the network contains the black links; for $\beta$ = 0.25, the network is formed by removing the marked black links and regrowing the green links). (b1) Link removal process. (b2) Formula for determining which links to remove. (c1) Removal of inactive neurons caused by link removal. (c2) Adjust and remove incomplete links caused by inactive neuron removal. (d1) Regrowth of links according to the CH2/3-L3 node-based soft rule. (d2) Detailed illustration of the CH2/3-L3 node-based soft rule. (e) Finished state of the network after one iteration. The next iteration repeats the steps (b1) - (e) from this finished state. $\tilde{A}$ indicates the removal set of the iteration and $A^{*}$ is the regrown set.
@@
To evaluate the effectiveness of the Cannistraci-Hebb Training soft rule with sigmoid density decay (CHTss), we conduct extensive experiments across multiple architectures and tasks. Firstly, to evaluate the basic concept of CHTs, we employ MLPs on benchmark datasets, including MNIST (LeCun et al., 1998), EMNIST (Cohen et al., 2017), and Fashion MNIST (Xiao et al., 2017). The results show that CHTs achieves superior performance compared to fully connected networks with only 1% of the connections (99% sparsity) in MLPs. Further, to evaluate the end-to-end approach CHTss, we utilize Transformers (Vaswani et al., 2017) on machine translation datasets such as Multi30k en-de (Elliott et al., 2016), IWSLT14 en-de (Cettolo et al., 2014), and WMT17 en-de (Bojar et al., 2017). Additionally, we evaluate CHTss also on the LLaMA-130M model with language modeling tasks and zero-shot evaluation tasks. From the results of the above experiments, CHTss outperforms fully connected Transformers with only 5% of the links on Multi30k and IWSLT and achieves performance comparable to the fully connected LLaMA-130M in language modeling tasks on OpenWebText. Moreover, CHTss outperforms the fully connected LLaMA-130M on zero-shot evaluation tasks on GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020) with only 30% density. These findings underscore the potential of CHTs and CHTss in enabling highly efficient and effective large-scale sparse neural network training.
@@ 2.1 Dynamic sparse training
Dynamic sparse training is a subset of sparse training methodologies. Unlike the static sparse training (also known as pruning at initialization) methods (Prabhu et al., 2018; Lee et al., 2019; Dao et al., 2022; Stewart et al., 2023), dynamic sparse training allows for the evolution of network topology during the training process. The pioneering method in this field was Sparse Evolutionary Training (SET) (Mocanu et al., 2018), which removes links based on the magnitude of their weights and regrows new links randomly. Subsequent developments have sought to refine and expand upon this concept of dynamic topological evolution. One such advancement was proposed by DeepR (Bellec et al., 2017), a method that adjusts network connections based on stochastic gradient updates combined with a Bayesian-inspired update rule. Another significant contribution is the RigL (Evci et al., 2020), which leverages the gradient information of non-existing links to guide the regrowth of new connections during training. MEST (Yuan et al., 2021) utilizes both gradient and weight magnitude information to selectively remove and randomly regrow new links, which is the same as SET. In addition, it introduces an EM&S strategy that allows the model training with a larger density and finally convergence to the desired density. The Top-KAST (Jayakumar et al., 2020) method maintains constant sparsity throughout training by selecting the top $K$ parameters based on parameter magnitude at each training step and applying gradients to a broader subset $B$, where $B\supset A$. To avoid settling on a suboptimal sparse subset, Top-KAST also introduces an auxiliary exploration loss that encourages ongoing adaptation of the mask. Additionally, sRigL (Lasby et al., 2023) adapts the principles of RigL to semi-structured sparsity, facilitating the training of vision models from scratch with actual speed-ups during training phases. Despite these advancements, the state-of-the-art method remains RigL-based, yet it is not fully sparse in backpropagation, necessitating the computation of gradients for non-existing links. Addressing this limitation, Zhang et al. (Zhang et al., 2024b) propose CHT, a dynamic sparse training methodology that adopts a gradient-free regrowth strategy that relies solely on topological information (network shape intelligence), achieving an ultra-sparse configuration that surpasses fully connected networks in some tasks.
@@ 2.2 Cannistraci-Hebb Theory and Network Shape Intelligence
As the SOTA gradient-free link regrown method, CHT (Zhang et al., 2024b) originates from a brain-inspired network science theory. Drawn from neurobiology, Hebbian learning was introduced in 1949 (Hebb, 1949) and can be summarized in the axiom: “neurons that fire together wire together.” This could be interpreted in two ways: changing the synaptic weights (weight plasticity) and changing the shape of synaptic connectivity (Cannistraci et al., 2013; Daminelli et al., 2015; Durán et al., 2017; Cannistraci, 2018; Narula, 2017). The latter is also called epitopological plasticity (Cannistraci et al., 2013) because plasticity means “to change shape,” and epitopological means “via a new topology.” Epitopological Learning (EL) (Daminelli et al., 2015; Durán et al., 2017; Cannistraci, 2018) is derived from this second interpretation of Hebbian learning and studies how to implement learning on networks by changing the shape of their connectivity structure. One way to implement EL is via link prediction, which predicts the existence and likelihood of each nonobserved link in a network. CH3-L3 is one of the best and most robust performing network automata which is inside Cannistraci-Hebb (CH) theory (Muscoloni et al., 2022) that can automatically evolve the network topology with the given structure. The rationale is that, in any complex network with local-community organization, the cohort of nodes tends to be co-activated (fire together) and to learn by forming new connections between them (wire together) because they are topologically isolated in the same local community (Muscoloni et al., 2022). This minimization of the external links induces a topological isolation of the local community, which is equivalent to forming a barrier around the local community. The external barrier is fundamental to maintaining and reinforcing the signaling in the local community, inducing the formation of new links that participate in epitopological learning and plasticity.
@@ 3 Cannistraci-Hebb training soft rule with sigmoid gradual density decay
3.1 Cannistraci-Hebb soft removal and regrowth
@@
Let $A_{t}$ be the set of existing links in the network at the training step $t$. Let $\tilde{A}_{t}$ be the set of removal links and $A_{t}^{*}$ be the set of regrown links. The overlap set between removed and regrown links at step $t$ can be quantified as $O_{t}=\tilde{A}_{t}\cap A_{t}^{*}$. An ELM occurs if the size of $O_{t}$ at step $t$ is significantly large compared to the size of $A_{t}^{*}$, indicating a high probability of the same links being removed and regrown repeatedly throughout the subsequent training steps. This can be formally represented as $\frac{|O_{t}|}{|A_{t}^{*}|}\geq\theta$, where $\theta$ is a predefined threshold close to 1, indicating strong overlap. This definition is essential for the understanding of CHT, as evidenced by the article (Zhang et al., 2024b) indicating that the overlap rate between removed and regrown links becomes significantly high within just a few epochs, leading to rapid topological convergence towards the ELM. Previously, CHT implements a topological early stop strategy to avoid predicting the same links iteratively. However, it will stop the topological exploration very fast and potentially trap the model within the ELM.
In this article, we adopt a probabilistic approach where the process of regrowth can be viewed as sampling from a $\{0,1\}$ multinomial distribution, with the score assigned by either removal metrics or link prediction scores, introducing a ”soft sampling” mechanism. In this setup, each mask value is not rigidly determined by the link prediction scores but allows for selecting (with lower probability) low-score links as new links, facilitating the escape from the epitopological local minima (ELM).
@@
As illustrated in Equation 1, RI assesses connections by normalizing the absolute weight of links that share the same input or output neurons. This method does not require calibration data and can perform comparably to the baseline post-training pruning methods like sparsegpt (Frantar & Alistarh, 2023) and wanda (Sun et al., 2023). Generally, WM and RI are straightforward, effective, and quick to implement in DST for link removal but give different directions for network percolation. WM prioritizes links with higher weight magnitudes, leading to rapid network percolation, whereas RI inherently values links connected to lower-degree nodes, thus maintaining a higher active neuron post-percolation (ANP) rate. The ANP rate is the ratio of the number of active neurons after training compared to the original number of neurons before training. These methods are equally valid but cater to different scenarios. For instance, using RI significantly improves results on the Fashion MNIST dataset compared to WM, whereas WM performs better on the MNIST and EMNIST datasets.
@@
Another significant challenge for CHT lies in the time complexity of link prediction. In the original CHT framework, the CH3-L3 metric is employed for link regrowth, defined as follows:

### v2

# Brain network science modelling of sparse neural networks enables Transformers and LLMs to perform as fully connected

@@ Abstract
This study aims to enlarge our current knowledge on the application of brain-inspired network science principles for training artificial neural networks (ANNs) with sparse connectivity. Dynamic sparse training (DST) emulates the synaptic turnover of real brain networks, reducing the computational demands of training and inference in ANNs. However, existing DST methods face difficulties in maintaining peak performance at high connectivity sparsity levels. The Cannistraci-Hebb training (CHT) is a brain-inspired method that is used in DST for growing synaptic connectivity in sparse neural networks. CHT leverages a gradient-free, topology-driven link regrowth mechanism, which has been shown to achieve ultra-sparse (1% connectivity or lower) advantage across various tasks compared to fully connected networks. Yet, CHT suffers two main drawbacks: (i) its time complexity is $\mathcal{O}(N\cdot d^{3})$- N node network size, d node degree - hence it can be efficiently applied only to ultra-sparse networks. (ii) it rigidly selects top link prediction scores, which is inappropriate for the early training epochs, when the network topology presents many unreliable connections. Here, we design the first brain-inspired network model - termed bipartite receptive field (BRF) - to initialize the connectivity of sparse artificial neural networks. Then, we propose a matrix multiplication GPU-friendly approximation of the CH link predictor, which reduces the computational complexity to $\mathcal{O}(N^{3})$, enabling a fast implementation of link prediction in large-scale models. Moreover, we introduce the Cannistraci-Hebb training soft rule (CHTs), which adopts a flexible strategy for sampling connections in both link removal and regrowth, balancing the exploration and exploitation of network topology. Additionally, we propose a sigmoid-based gradual density decay strategy, leading to an advanced framework referred to as CHTss. Empirical results show that BRF offers performance advantages over previous network science models. Using 1% of connections, CHTs outperforms fully connected networks in MLP architectures on visual classification tasks, compressing some networks to less than 30% of the nodes. Using 5% of the connections, CHTss outperforms fully connected networks in two Transformer-based machine translation tasks. Finally, using 30% of the connections, CHTs and CHTss achieve superior performance compared to other dynamic sparse training methods in language modeling across different sparsity levels, and it surpasses the fully connected counterpart in zero-shot evaluations.
@@ 1 Introduction
Artificial neural networks (ANNs) have led to significant advances in various fields such as natural language processing, computer vision, and deep reinforcement learning. The most common ANNs consist of several fully connected (FC) layers, which account for a large portion of the total parameters in recent large language models touvron2023llama ; zhang2022opt . This dense connectivity poses major challenges during the model training and deployment phases. In contrast, neural networks in the brain inherently exhibit sparse connectivity drachman2005we ; walsh2013peter . This natural design in the brain exploits sparsity, suggesting a model in which the number of connections does not scale quadratically with the number of neurons. This could alleviate computational constraints, enabling more scalable network architectures.
Dynamic sparse training (DST) SET ; jayakumar2020top ; RIGL ; MEST ; zhang2023ultra has emerged as a promising approach to reduce computational and memory overhead of training deep neural networks while maintaining or even improving model performance. DST is also biologically inspired: it draws an analogy to synaptic turnover holtmaat2009synaptic in the brain, a fundamental neurobiological process in which synapses are continuously formed, strengthened, weakened, and eliminated over time. This dynamic rewiring enables the brain to adapt, learn, and store memories efficiently while preserving the overall stability of the network. Similarly, in DST, connections are dynamically pruned and regrown throughout training, allowing the network to adapt its connectivity structure in response to learning signals while maintaining a fixed sparsity level.
@@
Apart from some detailed distinctions, the primary innovation in this field centers on the development of the regrowth criterion. A notable advancement is the gradient-free regrowth method introduced by Cannistraci-Hebb training (CHT) zhang2023ultra . This method is inspired by epitopological learning—literally meaning ‘new topology’—and is rooted in brain-inspired network science theory cannistraci2013link ; 18 ; 19 ; 20 ; Narula . Epitopological learning explores how learning can be implemented on complex networks by changing the shape of their connectivity structure (epitopological plasticity). CHT has demonstrated remarkable advantages in training ultra-sparse ANNs with connectivity levels of 1% or lower, often outperforming fully connected networks in various tasks. However, despite these advances, CHT encounters two major challenges: 1) During dynamic sparse training, its rigid link selection mechanism can lead to epitopological local minima where the sets of removed links and regrown links exhibit significant overlap, severely impedes the exploration of optimal network topologies. 2) The time complexity of the CHT regrowth method, Cannistraci-Hebb 3 on Length 3 paths (CH3-L3p), is $\mathcal{O}(N\cdot d^{3})$, where $N$ represents the number of nodes in the network and $d$ is the average degree. A length 3 path is a walk of three consecutive links. As the network becomes denser, the time complexity approaches $\mathcal{O}(N^{4})$, rendering it impractical for large-scale and higher-density models.
@@
In this article, we present the Cannistraci-Hebb Training soft rule (CHTs), which introduces several key innovations: 1) To address the issue of epitopological local minima, CHTs employs a multinomial distribution to sample link scores from removal and regrowth metrics, enabling more flexible and effective exploration of network topologies. 2) CHTs incorporates novel substitution node-based link prediction mechanisms, reducing the computational time complexity to $\mathcal{O}(N^{3})$. This improvement makes CHTs scalable to large-scale models with higher network density. 3) CHTs initializes the sparse topologies for bipartite networks with the brain-like receptive field, demonstrating superior performance compared to the traditional Erdős–Rényi SET , bipartite small world, and bipartite scale-free model zhang2023ultra . Additionally, we propose a sigmoid gradual density decay strategy, which, when integrated with CHTs, forms an enhanced framework termed CHTss. This combined approach further optimizes the training process for sparse neural networks.
@@
Figure 1: Illustration of the CHTs process. One training iteration follows the steps of (a1) $\rightarrow$ (b1) $\rightarrow$ (c1) $\rightarrow$ (c2) $\rightarrow$ (d1) $\rightarrow$ (e). (a1) Network initialization with each of the sandwich layers (bipartite networks connecting layers’ input nodes to their output nodes) being a bipartite receptive field (BRF) network. (a2) BRF network representation with $r$ = 0. (b1) Link removal process. (b2) Formula for determining which links to remove. (c1) Removal of inactive neurons caused by link removal. (c2) Adjust and remove incomplete links caused by inactive neuron removal. (d1) Regrowth of links according to the CH2-L3 node-based soft rule. (d2) Detailed illustration of the CH2-L3 node-based soft rule. (e) Finished state of the network after one iteration. The next iteration repeats the steps (b1) - (e) from this finished state. $\tilde{A}$ indicates the removal set of the iteration and $A^{*}$ is the regrown set.
@@
To evaluate the effectiveness of the Cannistraci-Hebb Training soft rule with sigmoid density decay (CHTss), we conduct extensive experiments across multiple architectures and tasks. Firstly, to assess the basic concept of CHTs, we employ MLPs on benchmark datasets, including MNIST MNIST , EMNIST cohen2017emnist , and Fashion MNIST xiao2017fashion . The results show that CHTs performs better than fully connected networks with only 1% of the connections (99% sparsity) in MLPs. Further, to evaluate the end-to-end approach CHTss, we utilize Transformers Transformer on machine translation datasets such as Multi30k en-de multi30k , IWSLT14 en-de cettolo-etal-2014-report , and WMT17 en-de bojar2017wmt . Additionally, we test CHTs and CHTss also on the LLaMA-130M model with language modeling tasks and zero-shot evaluation tasks. From the results of the above experiments, CHTss outperforms fully connected Transformers with only 5% of the links on Multi30k and IWSLT and achieves performance comparable to the fully connected LLaMA-130M in language modeling tasks on OpenWebText. Moreover, CHTs and CHTss outperform the fully connected LLaMA-130M on zero-shot evaluation tasks on GLUE wang2019glue and SuperGLUE wang2020supergluestickierbenchmarkgeneralpurpose with only 30% connections. These findings underscore the potential of CHTs and CHTss in enabling highly efficient and effective large-scale sparse neural network training.
@@ 2.1 Dynamic sparse training
Dynamic sparse training is a subset of sparse training methodologies. Unlike static sparse training methods (also known as pruning at initialization) prabhu2018deep ; SNIP ; dao2022monarch ; stewart2023data , dynamic sparse training allows for the evolution of network topology during the training process. The pioneering method in this field is Sparse Evolutionary Training (SET) SET , which removes links based on the magnitude of their weights and regrows new links randomly. Subsequent developments have sought to refine and expand upon this concept of dynamic topological evolution. One such advancement was proposed by DeepR bellec2017deep , a method that adjusts network connections based on stochastic gradient updates combined with a Bayesian-inspired update rule. Another significant contribution is RigL RIGL , which leverages the gradient information of non-existing links to guide the regrowth of new connections during training. MEST MEST utilizes both gradient and weight magnitude information to selectively remove and randomly regrow new links, analogously to SET. In addition, it introduces an EM&S strategy that allows the model to train at a higher density and gradually converge to the target sparsity. The Top-KAST jayakumar2020top method maintains constant sparsity throughout training by selecting the top $K$ parameters based on parameter magnitude at each training step and applying gradients to a broader subset $B$, where $B\supset A$. To avoid settling on a suboptimal sparse subset, Top-KAST also introduces an auxiliary exploration loss that encourages ongoing adaptation of the mask. Additionally, sRigL lasby2023srigl adapts the principles of RigL to semi-structured sparsity, facilitating the training of vision models from scratch with actual speed-ups during training phases. Despite these advancements, the state-of-the-art method remains RigL-based, yet it is not fully sparse in backpropagation, necessitating the computation of gradients for non-existing links. Addressing this limitation, Zhang et al. zhang2023ultra propose CHT, a dynamic sparse training methodology that adopts a gradient-free regrowth strategy that relies solely on topological information (network shape intelligence), achieving an ultra-sparse configuration that surpasses fully connected networks in some tasks.
@@ 2.2 Cannistraci-Hebb Theory and Network Shape Intelligence
As the SOTA gradient-free link regrown method, CHT zhang2023ultra originates from a brain-inspired network science theory. Drawn from neurobiology, Hebbian learning was introduced in 1949 (hebbian, ) and can be summarized in the axiom: “neurons that fire together wire together.” This could be interpreted in two ways: changing the synaptic weights (weight plasticity) and changing the shape of synaptic connectivity (cannistraci2013link, ; 18, ; 19, ; 20, ; Narula, ). The latter is also called epitopological plasticity (cannistraci2013link, ) because plasticity means “to change shape,” and epitopological means “via a new topology.” Epitopological Learning (EL) (18, ; 19, ; 20, ) is derived from this second interpretation of Hebbian learning and studies how to implement learning on networks by changing the shape of their connectivity structure. One way to implement EL is via link prediction, which predicts the existence and likelihood of each nonobserved link in a network. CH3-L3 is one of the best-performing and most robust network automata, belonging to the Cannistraci-Hebb (CH) theory (CHA, ), which can automatically evolve the network topology starting from a given structure. The rationale is that, in any complex network with local-community organization, the cohort of nodes tends to be co-activated (fire together) and to learn by forming new connections between them (wire together) because they are topologically isolated in the same local community (CHA, ). This minimization of the external links induces a topological isolation of the local community, which is equivalent to forming a barrier around it. The external barrier is fundamental to maintaining and reinforcing the signaling in the local community, inducing the formation of new links that participate in epitopological learning and plasticity.
@@ 3 Cannistraci-Hebb training soft rule with sigmoid gradual density decay
3.1 Soft removal and regrowth.
@@
Let $A_{t}$ be the set of existing links in the network at the training step $t$. Let $\tilde{A}_{t}$ be the set of removal links and $A_{t}^{*}$ be the set of regrown links. The overlap set between removed and regrown links at step $t$ can be quantified as $O_{t}=\tilde{A}_{t}\cap A_{t}^{*}$. An ELM occurs if the size of $O_{t}$ at step $t$ is significantly large compared to the size of $A_{t}^{*}$, indicating a high probability of the same links being removed and regrown repeatedly throughout the subsequent training steps. This can be formally represented as $\frac{|O_{t}|}{|A_{t}^{*}|}\geq\theta$, where $\theta$ is a predefined threshold close to 1, indicating strong overlap. This definition is essential for the understanding of CHT, as evidenced by the article zhang2023ultra indicating that the overlap rate between removed and regrown links becomes significantly high within just a few epochs, leading to rapid topological convergence towards the ELM. Previously, CHT implemented a topological early stop strategy to avoid predicting the same links iteratively. However, it will stop the topological exploration very fast and potentially trap the model within the ELM.
In this article, we adopt a probabilistic approach where the regrowth process is modeled as sampling from a $\{0,1\}$ multinomial distribution, with probabilities determined by link prediction scores, thereby introducing a "soft sampling" mechanism. Under this formulation, each mask value is not rigidly dictated by the link prediction score; instead, low-score links can still be selected with lower probability, facilitating escape from epitopological local minima (ELM). To demonstrate that soft sampling effectively balances exploration and exploitation, we evaluate its behavior in Figure 3, which presents the impact of varying softness levels during training of LLaMA-60M for 5000 steps under 90% sparsity.
We compare the in-time over-parameterization (ITOP) rate liu2021itop , which quantifies the cumulative proportion of links activated throughout training. As shown, deterministic regrowth leads to rapid topological convergence after approximately 1000 steps, indicating that it becomes trapped in an ELM without further exploration. Random regrowth, while capable of escaping ELMs by introducing new connections, fails to exploit the underlying topological structure effectively. In contrast, soft regrowth achieves a balance by both exploiting the current topology and exploring new link combinations probabilistically. This balance enables a more appropriate exploration of the connectivity space, ultimately leading to superior performance, as evidenced by the results.
@@
As illustrated in Equation 1, RI assesses connections by normalizing the absolute weight of links that share the same input or output neurons. This method does not require calibration data and can perform comparably to the baseline post-training pruning methods like sparsegpt frantar2023sparsegpt and wanda sun2023wanda . Generally, WM and RI are straightforward, effective, and quick to implement in DST for link removal, but give different directions for network percolation. WM prioritizes links with higher weight magnitudes, leading to rapid network percolation, whereas RI inherently values links connected to lower-degree nodes, thus maintaining a higher active neuron post-percolation (ANP) rate. The ANP rate is the ratio of the number of active neurons after training over the original number of neurons before training. These methods are equally valid but cater to different scenarios. For instance, using RI significantly improves results on the Fashion MNIST dataset compared to WM, whereas WM performs better on the MNIST and EMNIST datasets.
@@
Figure 2: One-time Link Prediction Runtime Performance Evaluation of node-based and path-based methods across varying densities and network sizes. In (a), the network size is fixed at 1024 × 1024, while in (b), the density is fixed at 5%.
@@
Another significant challenge for CHT lies in the time complexity of link prediction. In the original CHT framework zhang2023ultra , the path-based CH3-L3p metric is employed for link regrowth, as discussed in Appendix C. However, this method incurs a high computational cost due to the need to compute and store all length-three paths, resulting in a time complexity of $O(N\cdot d^{3})$, where $N$ is the number of nodes and $d$ is the network’s average degree. This complexity is prohibitive for large models with numerous nodes and higher-density layers.
To address this issue, we introduce a more efficient, node-based paradigm that eliminates the reliance on length-three paths between seed nodes. Instead, this approach focuses on the common neighbors of seed nodes. The node-based version of CH3-L3p, denoted as CH2-L3n, also depends on the internal local community links (iLCL, the links between the common neighbors CHA ) to enhance the expressiveness of the formula. This variant is formulated as: